OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

\[(7+3\sqrt{2)(9-5}\sqrt{2)}\] is the problem.

OpenStudy (anonymous):

The answer choices:

OpenStudy (anonymous):

A. \[33-8\sqrt{2}\] B. \[63-15\sqrt{2}\] C. \[27\sqrt{2}-35\]

hero (hero):